OpenStudy (anonymous):

you travel 10 mi on your bicycle in the same amount of time it takes to travel 8 mi on his bicycle. If your friend rides his bike 2mi/h slower than you ride your bike, FIND THE RATE at which each of you is traveling

jimthompson5910 (jim_thompson5910):

Let r = your speed, so your friend is going r-2 mph Equation for your speed/distance d = rt 10 = r*t r = 10/t ------------ Equation for your friend's speed and distance d = rt 8 = (r-2)t 8 = (10/t - 2)t 8 = 10 - 2t keep going to solve for t once you find t, use it in r = 10/t to find r